How would I solve " 11 divided by the fourth root of 11"? Thanks.
Answer by rothauserc(4718)

we write 11 divided by the fourth root of 11 as
11 / 11^1/4 = 11 * 11^-1/4 = 11^3/4
==================================================
11^3/4 is the 4th root of 11 cubed